OpenStudy (anonymous):

Object A has a mass of 100 grams. Object B has a mass of 150 grams. They are both traveling at the same velocity. Which can you conclude?

OpenStudy (anonymous):

have different kinetic energy! if the speed is 10 m / s the kinetic energy of the car would be: Ec = (100.10 ^ 2) / 2 Ec = 5000 the kinetic energy of the car B would be: Ec = (150.10 ^ 2) / 2 Ec = 7500